P=p(1+R)^2
i need to solve for R.

Answer by Boreal(15235)

Isolate the R
so P/p=(1+R)^2
take the square root of both sides
sqrt(P/p)=+/- (1+R)
1+R=sqrt(P/p)
R=sqrt(P/p)-1
and
-1-R= sqrt (P/p
R=-1-sqrt(P/p)